OpenStudy (2yoututu11):

Evaluate the expression by using the given value of the variable. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 A. 36 B. 21 C. 3 D. –3

OpenStudy (wolfboy):

Plug in the numbers 16/2*4-10/5+6 is the equation now

OpenStudy (wolfboy):

Now we do 16/2=8 8*4=32 now are equation is 32-10/5+6

OpenStudy (wolfboy):

BUT the order of operations PEMDAS we have to fallow them

OpenStudy (wolfboy):

10/5=2

OpenStudy (wolfboy):

now we have 32-2+6

OpenStudy (wolfboy):

32-2=30 30+6=36 our answer is a

OpenStudy (2yoututu11):

thanks

OpenStudy (wolfboy):

your welcome